OpenStudy (anonymous):

2a + 2b = c for b How do I solve for the indicated variable?

OpenStudy (jhannybean):

You want to isolate the b on one side and everything else on the other. Subtract -2a on both sides of equation. -2a + 2a + 2b = c - 2a Simplify 2b = - 2a + c divide by 2 on both sides of the equation. 2b/2 = ( -2a + c)/2 b= -2a/2 + c/2 Simplify once again. b = -a + c/2